OpenStudy (anonymous):

y''-8y'+12y=0, y(0)=4, y(1)=2

OpenStudy (anonymous):

So this is a homogenous differential equation. So what we want to do is take the characteristic polynomial and find the roots of that equation. Think you can do that part?

OpenStudy (anonymous):

yes i have done that part y=c1e^(2x)+c2e^(6x)

OpenStudy (anonymous):

Hmm....Did you change the equation or was I just not paying attention? O.o

OpenStudy (anonymous):

sorry i made mistake it was 12y not 8y

OpenStudy (anonymous):

no i changed it

OpenStudy (anonymous):

Ah, okay : ) Then yes, your complimentary solution is correct then. So the first of your initial conditions gives you y(0) = 4. So this means when x is 0, I need to be equal to 4 4 = c1e^(2*0)+c2e^(6*0). This of course simplifies to 4 = c1 + c2. So not enough information to solve for those coefficients. So we need to use the other initial condition. Did you mean that second condition to be y(1) or y'(1)?

OpenStudy (anonymous):

yes i have this first question same but second question is not y' because the second value is given is y not y'

OpenStudy (anonymous):

I just wanted to sure it was supposed to be y, you often see one condition with y and the other with y' Okay, so using the 2nd condition y(1) = 2, we have 2 = c1e^2 + c2e^6. So this gives us two equations which we can use to solve for c1 and c2. 4 = c1 + c2 2 = c1e^2 + c2e^6 So now its just using whichever method of solving a system of equations you prefer.

OpenStudy (anonymous):

i know we have two questions and i get c1=10.908 and c2= -14.908 but i am not sure if they are right

OpenStudy (anonymous):

Well, let's check. So if I were to just do substitution, I'll do 4-c2 = c1 2 = (4-c2)e^2 + c2e^6 2 = 4e^2 - c2e^2 + c2e^6 2-4e^2 = c2(-e^2 + e^6) \[c_{2}=\frac{ 2-4e^{2} }{ e^{6}-e^{2} }\approx -.07\] With that value, Ill plug it in for c2 into the first equation: 4 = c1 - .07 \(c_{1}\approx\ 4.07\)

OpenStudy (anonymous):

the way i did i found the answer first for e^2 and e^6 for second question and put it back to first question

OpenStudy (anonymous):

no when i put c1 as 4.07 and c2=-0.07 i didnt get right it

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

well, those are rounded off answers. More decimals to those answers would give: \(c_{1} \approx4.06958\) \(c_{2} \approx-.069579\) Is there any requirement on rounding or if it wants an exact solution or not?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Another possible approach is using the Laplace transform, if you're familiar with it.